OpenStudy (anonymous):

y=square root of x-4, inverse of the function

OpenStudy (anonymous):

x(y)=y^2+8y+16

OpenStudy (bahrom7893):

is it: y=sqrt(x)-4?

OpenStudy (bahrom7893):

or is it y=sqrt(x-4)

OpenStudy (anonymous):

(x-4)

OpenStudy (bahrom7893):

y=sqrt(x-4) solve for x: y^2=x-4 x=y^2+4. Now just swap x and y: y=x^2+4 is the inverse

OpenStudy (campbell_st):

the inverse is y = x^2 + 4